### v1

@@ 4.3.3 Finite-dimensionalizing linear inverse problems
One exception where analysis can be pushed through is the method of data-driven regularization-by-projection [224, 225]. The analysis adopts the framework of functional-analytic regularization [233] and works in the setting that the forward operator is unknown. Instead, the forward map is only accessible through input-output training data. The theory establishes convergence guarantees in the limit of infinite training data, both for noise-free and noisy observations. We now survey such results.
@@
The motivation for the formula (4.30) is that the action of the forward map $\mathcal{A}$ on the span $\mathcal{U}_{N}$ of the input training data is completely determined by the data. It is then possible to deduce that $\mathcal{R}_{N}y$ corresponds to the minimum norm solution of the input-projected linear operator equation $y=\mathcal{A}P_{\mathcal{U}_{N}}u$. Then $\mathcal{R}_{N}$ in (4.30) is seen to be the Moore–Penrose pseudoinverse of $\mathcal{A}P_{\mathcal{U}_{N}}$ [225, Thm. 4, p. 5].
@@
are nested and that $\overline{\cup_{N\in\mathbb{N}}\mathcal{U}_{N}}=\mathcal{U}$. Notice that the training data need not be generated according to a probability distribution; they could be deterministic and highly correlated. For example, one could select the $\{u_{n}\}$ to be a problem-adapted orthonormal basis or construct them according to some greedy algorithm. Under further technical assumptions on the compatibility between $\mathcal{A}$, the linear span $\mathcal{U}_{N}$, and a fixed $y\in\operatorname{Ran}(\mathcal{A})$ [224, Sec. 2.2.1], it holds that $\mathcal{R}_{N}y$ converges both weakly and strongly to $\mathcal{A}^{-1}y$ in the Hilbert space $\mathcal{U}$ as $N\to\infty$ [225, Thms. 9, 11, 15]. We can also handle the noisy data case $y^{\delta}\not\in\operatorname{Ran}(\mathcal{A})$. Optimizing a standard bias-variance tradeoff [234, Thm. 46, p. 33] by choosing $N=N(\delta)\to\infty$ according to the requirement
@@
by [225, Thm. 17, p. 17]. In (4.36), we write $B_{\delta}(y)\coloneqq{\mathopen{}\mathclose{{\left\{z\in\mathcal{Y}\,\middle|\,\lVert y-z\rVert\leq\delta}}\right\}}$.
The convergence result (4.36) is pointwise in the “input space” $\operatorname{Ran}(\mathcal{A})$ and uniform over perturbations. Because $\mathcal{A}^{-1}$ is not continuous, we cannot expect a version of (4.36) that is uniform over the unit ball of $\operatorname{Ran}(\mathcal{A})$ to be valid [234, Thm. 4.9, p. 36]. However, comparing to Eqs. 4.22, 4.24 and 4.26 which are uniform over the domain of a nonlinear inverse map, it is of interest to determine meaningful compact source sets or probability measures over parameter space that lead to uniform or average case convergence, respectively, of $\mathcal{R}_{N(\delta)}$ at the level of operators, possibly with respect to new accuracy metrics.
The data-driven regularization-by-projection method can be viewed both as learning the forward operator—via $\mathcal{A}\approx\mathcal{A}P_{\mathcal{U}_{N}}$—or as learning the inverse operator—via $\mathcal{A}^{-1}\approx\mathcal{A}^{-1}P_{\mathcal{Y}_{N}}$. Both viewpoints can be unified in the special case that the training data arise from the SVD of $\mathcal{A}$ [225, Remark 6, p. 6]. The current approach is but one of many ways to finite-dimensionalize linear inverse problems by discretizing parameters or data [235]. Dual least squares and explicitly regularized variational problems also enjoy convergence guarantees [225, Secs. 4–5]. Beyond deterministic data with span-increasing and linear independence constraints, other choices of projection sets could include the leading eigenspaces of convenient “prior” operators that are compatible with $\mathcal{A}$ [177]. This could help circumvent the technical difficulties associated with the Gram–Schmidt procedure.
@@ 4.4 Beyond point estimation: Data-to-posterior maps
Compared to data-to-parameter map learning, the literature on learning data-to-posterior maps in the context of infinite-dimensional Bayesian inverse problems (Section 2.1.1) is much sparser [34, Chp. 6]. For a fixed observation $y$, there is abundant work on how to compute the posterior probability measure $\mu^{y}$ given by (2.4), e.g., by sampling or density estimation methods. However, such traditional inference methods are relatively slow. This cost is magnified if the application requires the repeated calculation of the posterior for many different measurements, which is the case in sequential data assimilation [236] or Bayesian optimal experimental design [106]. By data-to-posterior operator learning, we refer to the problem of approximating the map $y\mapsto\mu^{y}$. This approach effectively solves entire families of Bayesian inverse problems parametrized by the observed data without the need for re-training or repeated optimization solves. Although the upfront training or data generation cost may be high, this offline cost is amortized if the trained map is evaluated multiple times and each evaluation is fast (e.g., near real-time). Similar considerations hold for the prior-to-posterior map $\mu\mapsto\mu^{y}$, or even the combined Bayesian solution map $(\mu,y)\mapsto\mu^{y}$, although these are outside the scope of the present chapter.
@@
To this end, one approach is to learn data-to-posterior maps with a hypernetwork [241]. The hypernetwork $\Psi$ maps the data $y$ to the parameters $\theta=\Psi(y)$ of another neural network $\mathsf{F}$. We view $\mathsf{F}$ as an approximate transport map from the prior $\mu$ to the posterior $\mu^{y}$. That is, $\mu^{y}\approx\mathsf{F}({\,\cdot\,};\theta)_{\#}\mu$. The map $\mathsf{F}$ only depends on $y$ through its parameters $\theta=\Psi(y)$. The training of $\Psi$ can be performed in an unsupervised way by optimizing the Kullback–Leibler (KL) divergence, but this requires knowledge of the likelihood to evaluate the loss. The overall method is a “meta” conditional normalizing flow and has been applied to inverse problems on function spaces [241]. Since the hypernetwork training task is difficult in practice, one should later fine tune the meta-conditional model or use it as a preconditioner for a fixed data realization if higher accuracy is required.
An alternative methodology is that of triangular transport maps [119]. This approach falls under the broader purview of computational measure transport [79], as described in Section 2.1.2. In (2.5), which we recall is the minimization problem $\min_{\mathsf{T}\in\mathcal{T}}\mathsf{d}(\nu,\mathsf{T}_{\#}\rho)$, the triangular framework takes $\mathsf{d}$ to be the KL divergence and $\mathcal{T}$ to be a class of triangular maps with a particular parametrization (e.g., polynomial or wavelet expansions) and structural constraints such as monotonicity [118]. This construction is inherently finite-dimensional, which precludes its scalability to infinite-dimensional inverse problems. One solution involves a relaxation to block triangular transport maps, which are well-defined between infinite-dimensional Banach spaces and naturally amortize posterior sampling with respect to the observation data [82]. Our discussion is restricted to a useful but specific case of a more general setup [81].
@@
for $(y,u)\in\mathcal{Y}\times\mathcal{U}$ and $\Psi\colon\mathcal{Y}\times\mathcal{U}\to\mathcal{U}$. The map $\Psi$ parametrizes the class $\mathcal{T}$. The nonlinear function (4.37) is called block triangular due to the variable dependence in each block coordinate that ensures the derivative of $\mathsf{T}$ takes values as a block lower-triangular linear operator. We further take $\rho\coloneqq\nu_{\mathcal{Y}}\otimes\rho_{\mathcal{U}}\in\mathscr{P}(\mathcal{Y}\times\mathcal{U})$ to be an independent product reference measure, where $\nu_{\mathcal{Y}}$ is the $\mathcal{Y}$-marginal of $\nu$ and $\rho_{\mathcal{U}}$ is arbitrary. A typical choice is $\rho_{\mathcal{U}}=\mu$ for a Bayesian inverse problem with prior $\mu$ [81, Sec. 4]. The form (4.37) is useful for the following reason. If $\mathsf{T}_{\#}\rho=\nu$, then
@@
for $\nu_{\mathcal{Y}}$-a.e. $y\in\mathcal{Y}$ [82, Thm. 2.4, p. 875]. That is, we can extract the desired conditional distribution $\nu({\,\cdot\,}\,|\,y)$ from the subcomponent $\Psi$ of the full transport map $\mathsf{T}$. Moreover, global minimizers $\mathsf{T}^{\star}$ to (2.5) over the class (4.37) exist and achieve $\mathsf{T}^{\star}\rho=\nu$ [82, Remark 2.11, p. 878].
These theoretical underpinnings help motivate the monotone generative adversarial network (M-GAN) architecture [82, Sec. 3], which is a powerful amortized solver for Bayesian inverse problems. M-GAN takes $\mathsf{d}$ to be a min-max approximation of the Wasserstein-$1$ distance. For PDE-based inverse problems such as Darcy flow, the transport map $\Psi$ is parametrized using PCA-Net (3.10). Practically, we only have access to $\nu$ through a finite training dataset of $N$ pairs of i.i.d. joint samples. Then $\nu$ in (2.5) is replaced by the empirical measure $\nu^{N}$ corresponding to these samples. To avoid overfitting to the training data, the reference $\rho$ is also replaced by its $M$-sample approximation, with $M$ possibly different from $N$. After training, the M-GAN estimator $\widehat{\Psi}$ satisfies
@@ 5 Learning prior distributions and regularizers
Learning prior information or regularization terms directly from data has become a powerful alternative to hand-crafted regularization in solving inverse problems [244, 245, 246, 137, 247, 248, 249, 250, 251]. Traditionally, priors were imposed using explicit penalty terms or probabilistic assumptions; however, modern approaches instead learn these priors from examples of signals or measurements. This allows one to capture complex structures or statistical correlations in data that are difficult to express analytically, especially in high-dimensional or ill-posed settings. The learned prior can be used either in variational inference or to guide iterative algorithms. This section focuses on two complementary paradigms for learning prior distributions and regularizers: approaches based on Bayesian variational inference in Section 5.1 and denoising-based regularization (e.g., plug-and-play methods) in Section 5.2.
5.1 Prior learning and Bayesian variational inference
In the Bayesian formulation, solving an inverse problem involves specifying a prior distribution $\mu$ over the unknown parameter $u$ and a likelihood function $\mathsf{L}^{y}(u)=\exp(-\Phi(u;y))$, which models the relationship between the data $y$ and the parameter $u$ (Section 2.1.1). The likelihood represents the forward and measurement processes as well as uncertainties, while the prior encodes our assumptions or background knowledge about $u$ before observing the data.
@@ 5.1 Prior learning and Bayesian variational inference
From this optimization perspective, regularization can be interpreted as imposing a prior through the term $R(u)\coloneqq-\log\mu(u)$. In classical inverse problems, handcrafted regularizers such as quadratic penalties for smoothness or sparsity-promoting norms play the role of $R$. For example, Gaussian priors lead to Tikhonov-type regularizers, while a Laplace prior aligns with sparsity-promoting regularizers. Note that in infinite dimensions, the densities appearing in the preceding formulas (5.1), (5.2), and (5.3) require careful interpretation [60].
The fact that the prior acts like a regularization term can also be seen from the variational form of Bayes’ theorem [34, 252]. Rather than viewing Bayes’ rule as a formula for conditioning, one can interpret the posterior $\mu^{y}$ as the minimizer of a variational problem over the metric space of probability measures. Specifically, the posterior distribution minimizes the functional
@@
where $\mathsf{KL}(\nu\;\|\;\mu)$ is the KL divergence between $\nu$ and the prior $\mu$. In this formulation, the KL divergence serves as a regularization term that penalizes deviations from the prior, while the expected negative log-likelihood promotes fidelity with the observed data. The choice of prior $\mu$ directly determines the structure and strength of the regularization term $\mathsf{KL}(\nu\;\|\;\mu)$.
In modern settings, data-driven approaches aim to learn the prior $\mu$ directly from training samples. By modeling complex, high-dimensional distributions of real-world data, these methods offer greater flexibility compared to traditional handcrafted priors. Machine learning techniques, particularly deep generative models [253], have been particularly effective in capturing such distributions. Examples include energy-based models, generative adversarial networks (GANs), variational autoencoders (VAEs), and score-based diffusion models [140, 254, 255, 256, 257, 258]. These learned priors provide a data-driven means to enforce regularization in the Bayesian framework and provide state-of-the-art reconstructions in many inverse problems.

### v2

@@ 4.3.3 Finite-dimensionalizing linear inverse problems
One exception where analysis can be pushed through is the method of data-driven regularization-by-projection [247, 248]. The analysis adopts the framework of functional-analytic regularization [256] and works in the setting that the forward operator is unknown. Specifically, the forward map is only accessible through input-output training data. The theory establishes convergence guarantees in the limit of infinite training data, both for noise-free and noisy observations. We now survey such results.
@@
The motivation for the formula (4.30) is that the action of the forward map $\mathcal{A}$ on the span $\mathcal{U}_{N}$ of the input training data is completely determined by the data. It is then possible to deduce that $\mathcal{R}_{N}y$ corresponds to the minimum norm solution of the input-projected linear operator equation $y=\mathcal{A}P_{\mathcal{U}_{N}}u$. Then $\mathcal{R}_{N}$ in (4.30) is seen to be the Moore–Penrose pseudoinverse of $\mathcal{A}P_{\mathcal{U}_{N}}$ [248, Thm. 4, p. 5].
@@
are nested and that $\overline{\cup_{N\in\mathbb{N}}\mathcal{U}_{N}}=\mathcal{U}$. Notice that the training data need not be generated according to a probability distribution; they could be deterministic and highly correlated. For example, one could select the $\{u_{n}\}$ to be a problem-adapted orthonormal basis or construct them according to some greedy algorithm. Under further technical assumptions on the compatibility between $\mathcal{A}$, the linear span $\mathcal{U}_{N}$, and a fixed $y\in\operatorname{Ran}(\mathcal{A})$ [247, Sec. 2.2.1], it holds that $\mathcal{R}_{N}y$ converges both weakly and strongly to $\mathcal{A}^{-1}y$ in the Hilbert space $\mathcal{U}$ as $N\to\infty$ [248, Thms. 9, 11, 15]. We can also handle the noisy data case $y^{\delta}\not\in\operatorname{Ran}(\mathcal{A})$. Optimizing a standard bias-variance tradeoff [257, Thm. 46, p. 33] by choosing $N=N(\delta)\to\infty$ according to the requirement
@@
by [248, Thm. 17, p. 17]. In (4.36), we write $B_{\delta}(y)\coloneqq{\mathopen{}\mathclose{{\left\{z\in\mathcal{Y}\,\middle|\,\lVert y-z\rVert_{\mathcal{Y}}\leq\delta}}\right\}}$.
The convergence result (4.36) is pointwise in the “input space” $\operatorname{Ran}(\mathcal{A})$ and uniform over perturbations. Because $\mathcal{A}^{-1}$ is not continuous, we cannot expect a version of (4.36) that is uniform over the unit ball of $\operatorname{Ran}(\mathcal{A})$ to be valid [257, Thm. 4.9, p. 36]. However, comparing to Eqs. 4.22, 4.24 and 4.26 which are uniform over the domain of a nonlinear inverse map, it is of interest to determine meaningful compact source sets or probability measures over parameter space that lead to uniform or average case convergence, respectively, of $\mathcal{R}_{N(\delta)}$ at the level of operators, possibly with respect to new accuracy metrics.
The data-driven regularization-by-projection method can be viewed both as learning the forward operator—via $\mathcal{A}\approx\mathcal{A}P_{\mathcal{U}_{N}}$—or as learning the inverse operator—via $\mathcal{A}^{-1}\approx\mathcal{A}^{-1}P_{\mathcal{Y}_{N}}$. Both viewpoints can be unified in the special case that the training data arise from the SVD of $\mathcal{A}$ [248, Remark 6, p. 6]. The current approach is but one of many ways to finite-dimensionalize linear inverse problems by discretizing parameters or data [258]. Dual least squares and explicitly regularized variational problems also enjoy convergence guarantees [248, Secs. 4–5]. Beyond deterministic data with span-increasing and linear independence constraints, other choices of projection sets could include the leading eigenspaces of convenient “prior” operators that are compatible with $\mathcal{A}$ [187]. This could help circumvent the technical difficulties associated with the Gram–Schmidt procedure.
@@ 4.4 Beyond point estimation: Data-to-posterior maps
Compared to data-to-parameter map learning, the literature on learning data-to-posterior maps in the context of infinite-dimensional Bayesian inverse problems (Section 2.1.1) is much sparser [36, Chp. 6]. For a fixed observation $y$, there is abundant work on how to compute the posterior probability measure $\mu^{y}$ given by (2.5), e.g., by sampling or density estimation methods. However, such traditional inference methods are relatively slow. This cost is magnified if the application requires the repeated calculation of the posterior for many different measurements, which is the case in sequential data assimilation [259] or Bayesian optimal experimental design [113]. By data-to-posterior operator learning, we refer to the problem of approximating the map $y\mapsto\mu^{y}$. This approach effectively solves entire families of Bayesian inverse problems parametrized by the observed data without the need for re-training or repeated optimization solves. Although the upfront training or data generation cost may be high, this offline cost is amortized if the trained map is evaluated multiple times and each evaluation is fast (e.g., near real-time). Similar considerations hold for the prior-to-posterior map $\mu\mapsto\mu^{y}$, or even the combined Bayesian solution map $(\mu,y)\mapsto\mu^{y}$, although these are outside the scope of the present chapter.
@@
To this end, one approach is to learn data-to-posterior maps with a hypernetwork [264]. The hypernetwork $\Psi$ maps the data $y$ to the parameters $\theta=\Psi(y)$ of another neural network $\mathsf{F}$. We view $\mathsf{F}$ as an approximate transport map from the prior $\mu$ to the posterior $\mu^{y}$. That is, $\mu^{y}\approx\mathsf{F}({\,\cdot\,};\theta)_{\#}\mu$. The map $\mathsf{F}$ only depends on $y$ through its parameters $\theta=\Psi(y)$. The training of $\Psi$ can be performed in an unsupervised way by optimizing the Kullback–Leibler (KL) divergence, but this requires knowledge of the likelihood to evaluate the loss. The overall method is a “meta” conditional normalizing flow and has been applied to inverse problems on function spaces [264]. Since the hypernetwork training task is difficult in practice, one should later fine tune the meta-conditional model or use it as a preconditioner for a fixed data realization if higher accuracy is required. Section 5.1.1 contains related discussion.
An alternative methodology is that of triangular transport maps [129]. This approach falls under the broader purview of computational measure transport [83], as described in Section 2.1.2. In (2.6), which we recall is the minimization problem $\min_{\mathsf{T}\in\mathcal{T}}\mathsf{d}(\nu,\mathsf{T}_{\#}\rho)$, the triangular framework takes $\mathsf{d}$ to be the Kullback–Leibler (KL) divergence and $\mathcal{T}$ to be a class of triangular maps with a particular parametrization (e.g., polynomial or wavelet expansions) and structural constraints such as monotonicity [128]. This construction is inherently finite-dimensional, which precludes its scalability to infinite-dimensional inverse problems. One solution involves a relaxation to block triangular transport maps, which are well-defined between infinite-dimensional Banach spaces and naturally amortize posterior sampling with respect to the observation data [86]. Our discussion is restricted to a useful but specific case of a more general setup [85].
@@
for $(y,u)\in\mathcal{Y}\times\mathcal{U}$ and $\Psi\colon\mathcal{Y}\times\mathcal{U}\to\mathcal{U}$. The map $\Psi$ parametrizes the class $\mathcal{T}$. The nonlinear function (4.37) is called block triangular due to the variable dependence in each block coordinate that ensures the derivative of $\mathsf{T}$ takes values as a block lower-triangular linear operator. We further take $\rho\coloneqq\nu_{\mathcal{Y}}\otimes\rho_{\mathcal{U}}\in\mathscr{P}(\mathcal{Y}\times\mathcal{U})$ to be an independent product reference measure, where $\nu_{\mathcal{Y}}$ is the $\mathcal{Y}$-marginal of $\nu$ and $\rho_{\mathcal{U}}$ is arbitrary. A typical choice is $\rho_{\mathcal{U}}=\mu$ for a Bayesian inverse problem with prior $\mu$ [85, Sec. 4]. The form (4.37) is useful for the following reason. If $\mathsf{T}_{\#}\rho=\nu$, then
@@
for $\nu_{\mathcal{Y}}$-a.e. $y\in\mathcal{Y}$ [86, Thm. 2.4, p. 875]. That is, we can extract the desired conditional distribution $\nu({\,\cdot\,}\,|\,y)$ from the subcomponent $\Psi$ of the full transport map $\mathsf{T}$. Moreover, global minimizers $\mathsf{T}^{\star}$ to (2.6) over the class (4.37) exist and achieve $\mathsf{T}^{\star}\rho=\nu$ [86, Remark 2.11, p. 878].
These theoretical underpinnings help motivate the monotone generative adversarial network (M-GAN) architecture [86, Sec. 3], which is a powerful amortized solver for Bayesian inverse problems. M-GAN takes $\mathsf{d}$ to be a min-max approximation of the Wasserstein-$1$ distance. For PDE-based inverse problems such as Darcy flow, the transport map $\Psi$ is parametrized using PCA-Net (3.10). Practically, we only have access to $\nu$ through a finite training dataset of $N$ pairs of i.i.d. joint samples. Then $\nu$ in (2.6) is replaced by the empirical measure $\nu^{N}$ corresponding to these samples. To avoid overfitting to the training data, the reference $\rho$ is also replaced by its $M$-sample approximation, with $M$ possibly different from $N$. After training, the M-GAN estimator $\widehat{\Psi}$ satisfies
@@ 5 Learning prior distributions and regularizers
Learning prior information or regularization terms directly from data has become a powerful alternative to hand-crafted regularization in solving inverse problems [267, 268, 269, 147, 270, 271, 272, 273, 274]. Traditionally, priors were imposed using explicit penalty terms or probabilistic assumptions; however, modern approaches instead learn these priors from examples of signals or measurements. This allows one to capture complex structures or statistical correlations in data that are difficult to express analytically, especially in high-dimensional or ill-posed settings. The learned prior can be used either in Bayesian inference or to guide iterative algorithms. This section focuses on two complementary paradigms for learning prior distributions and regularizers: approaches based on empirical Bayes-type prior learning in Section 5.1 and denoising-based regularization (e.g., plug-and-play methods) in Section 5.2.
5.1 Prior distribution learning
In the Bayesian formulation, solving an inverse problem involves specifying a prior distribution $\mu$ over the unknown parameter $u$ and a likelihood function $\mathsf{L}^{y}(u)=\exp(-\Phi(u;y))$ that models the relationship between the data $y$ and the parameter $u$ (Section 2.1.1). The likelihood represents the forward and measurement processes as well as uncertainties, while the prior encodes our assumptions or background knowledge about $u$ before observing the data.
@@ 5.1 Prior distribution learning
From this optimization perspective, regularization can be interpreted as imposing a prior through the term $R(u)\coloneqq-\log\mu(u)$. In classical inverse problems, handcrafted regularizers such as quadratic penalties for smoothness or sparsity-promoting norms play the role of $R$. For example, Gaussian priors lead to Tikhonov-type regularizers, while a Laplace prior aligns with sparsity-promoting regularizers.
Note that in infinite dimensions, the densities appearing in the preceding formulas (5.1), (5.2), and (5.3) require careful interpretation [64]. For certain Bayesian inverse problems on Banach spaces, an infinite-dimensional analog of Bayes’ rule holds (recall Section 2.1.1 and Eq. 2.5). This permits an analysis similar to that of Eqs. 5.1, 5.2 and 5.3, but remains valid in infinite dimensions. Although such extensions cannot rely on probability densities, natural generalizations of MAP estimators (5.2) exist based on small-ball probabilities; see [275, 276].
The fact that the prior acts like a regularization term can also be seen from the variational form of Bayes’ theorem [36, 277]. Rather than viewing Bayes’ rule as a formula for conditioning, one can interpret the posterior $\mu^{y}$ as the minimizer of a variational problem over the metric space of probability measures. Specifically, the posterior distribution satisfies
@@
where $\mathsf{KL}(\nu\;\|\;\mu)$ is the KL divergence between $\nu$ and the prior $\mu$. Equation (5.4) requires the absolute continuity of the posterior with respect to the prior and follows from the basic properties of KL divergence. In this formulation, the KL divergence serves as a regularization term that penalizes deviations from the prior, while the expected negative log-likelihood promotes fidelity with the observed data. The choice of prior $\mu$ directly determines the structure and strength of the regularization term $\mathsf{KL}(\nu\;\|\;\mu)$.
In modern settings, data-driven approaches aim to learn the prior $\mu$ directly from training samples. Most of these approaches fall under the empirical Bayes classification [278, 279] in which the prior $\mu$ is not modeled as random itself, unlike hierarchical Bayesian methods [280]. By modeling complex, high-dimensional distributions of real-world data, prior learning methods offer greater flexibility compared to traditional handcrafted priors. Machine learning techniques, particularly deep generative models [281], have been particularly effective in capturing such distributions. Examples include energy-based models, generative adversarial networks, variational autoencoders, and score-based diffusion models [150, 282, 283, 284, 285, 286]. These learned priors provide a data-driven means to enforce regularization in the Bayesian framework and provide state-of-the-art reconstructions in many inverse problems.